3. A certain car is capable of accelerating at a rate of 0.60 m/s^2. How long does it take for this car to go from a speed of 55 mi/h to a speed of 60 mi/h?
1
Expert's answer
2018-08-07T15:43:08-0400
The problem involves two different systems of units. We need to either convert "m/s2" to "mi/hr2" or convert mi/h to m/s.
Dimensional Analysis Formula:
1 mi/hr* (1mi/s)/(3600 mi/hr) *( 5280 ft/s)/(1 mi/s)*(1 m/s)/(3.2808 ft/s )=0.44704 m/s

So, the initial speed
v_i=55 mi/h=55×0.44704=24.59 m/s
The final speed
v_f=60 mi/h=60×0.44704=26.82 m/s

Since the acceleration of the car is constant, the relation between velocity and time is givenby the following formula:
v_f=v_i+at
Hence
t=(v_f-v_i)/a
Substituting values, we obtain
t=(26.82-24.59)/0.60=3.72 s